# Convergence for Natural Policy Gradient on Infinite-State Average-Reward Markov Decision Processes

### v1

@@ 5.4. Step 3: Bounding relative value $V_{k}(s)$ for all states $s$
$\forall s,k,\quad$ $V_{k}(s)\geq V_{0}(s)\frac{J_{*}-z}{J_{0}-z}-\tau^{bound}_{J_{0},z}\frac{(J_{*}-J_{0})(c_{\max}-z)(J_{*}-z)}{(J_{0}-z)^{2}}$
@@
We can upper bound each of these quantities. Note that $\bar{r}_{k}^{out}(s)\leq z$, and $\bar{r}_{k}^{in}(s)\leq c_{\max}$. Together, we can upper bound $V_{k}(s)$ relative to $\tau_{k}(s)$, or equivalently upper bound $\tau_{k}(s)$ relative to $V_{k}(s)$:
@@
Note that $-J_{k}+z<0$, so the direction of the inequality flips.
@@
$V_{k+1}(s)$ $\geq V_{k}(s)\frac{J_{k+1}-z}{J_{k}-z}+\tau^{bound}_{J_{0},z}(J_{k+1}-J_{k})\frac{(c_{\max}-z)}{J_{k}-z}$
@@
$V_{k}(s)$ $\geq V_{0}(s)\frac{J_{*}-z}{J_{0}-z}-\tau^{bound}_{J_{0},z}\frac{(J_{*}-J_{0})(c_{\max}-z)(J_{*}-z)}{(J_{0}-z)^{2}}$
@@
5.5. Step 4: Proof of main result
@@
Note that for any specific state $s$, $\sum_{k=1}^{T}Q_{k}(s,\pi_{*})-V_{k}(s)$ is exactly the difference in total reward between a specific fixed policy $\pi_{*}(a\mid s)$ and the weighted majority policy $\pi_{k}(a\mid s)$ of the expert advice problem with with reward function at time step $k$ of $Q_{k}(s,a)$.
@@
$J_{*}-J_{T}$ $\leq\frac{c_{*}}{\sqrt{T}}.\qed$
6. Queueing MDPs
We have proven Theorem 1, which gives a convergence rate bound for the Natural Policy Gradient (NPG) algorithm in infinite-state, average-reward MDPs that satisfy Assumptions 1 and 2, given an assumption on the quality of the initial policy. In this section, we demonstrate that a broad, natural family of MDPs arising from queueing theory, known as the generalized switch with static environment (GSSE) setting “(Stolyar, 2004), satisfy these assumptions. Moreover, we prove that a natural and well-studied policy, the MaxWeight policy, satisfies the requirement to be used as an initial policy in Theorem 1. We therefore demonstrate that NPG efficiently converges to the optimal policy in this broad natural class, when initialized with the MaxWeight policy.
We define the GSSE setting in Section 6.1. This setting is broad enough to capture a wide variety of natural queueing problems, including the N-system, the input-queued switch, and the multiserver-job setting. We primarily focus on the optimization goal of minimizing mean queue length $E[q]$, which corresponds to the reward function equal to the negative total queue length. Our results also apply to generalizations such as expected weighted queue length $E[cq]$ or the $\alpha$th moment of queue length $E[q^{\alpha}]$.
In Section 6.2, we discuss prior work on MaxWeight and the Generalized Switch setting. In Section 6.5, we prove that NPG with MaxWeight initialization meets the requirements for Theorem 1 and converges rapidly, using lemmas built up in Sections 6.3 and 6.4.
6.1. Definition of Generalized Switch with Static Environment
The Generalized Switch with Static Environment (GSSE) is a queueing model with $n$ classes of jobs, and the state of the system is a size-$n$ vector $q=\{q_{i}\}$ specifying the number of jobs of each class present in the system. This model is a special-case of the Generalized Switch model (Stolyar, 2004).
There are $m$ service options. At each time step, prior to new jobs arriving, the scheduling policy selects a service option $j$. If the policy selects service option $j$, then $w_{i}^{j}$ jobs of class $i$ complete, where $w_{i}^{j}$ is a bounded, i.i.d. random variable. Jobs of each class complete independently. If on some time step $t$, $w_{i}^{j}(t)$ exceeds $q_{i}(t)$, the number of jobs of class $i$ present in the system, then all of the class $i$ jobs complete.
Our primary reward function $r(s,\cdot)$ of state $s=\{q_{i}\}$ under any action is the negative total queue length $r(s,\cdot)=-\sum_{i}q_{i}$. We will also consider the setting where the reward is the negative $\alpha$-th moment of the total queue length, $-(\sum_{i}q_{i})^{\alpha}$, for a generic $\alpha\geq 1$.
6.2. Prior work on MaxWeight and Generalized Switch
The Generalized Switch is an expansive queueing model, simultaneously generalizing the input-queued switch and the parallel server system (Stolyar, 2004). An important scheduling policy in the Generalized Switch model is the MaxWeight policy (Tassiulas and Ephremides, 1990), which chooses the service option which maximizes the inner product of the queue lengths and the service rates:
(6) $\text{MaxWeight}(q):=\arg\max_{j}\sum\nolimits_{i}q_{i}\mu^{j}_{i}$
In particular, MaxWeight is known to be throughput optimal, meaning that it keeps the system stable for all arrival rates within the capacity region (Stolyar, 2004; Hurtado Lange and Maguluri, 2022). Moreover, MaxWeight asymptotically minimizes the total queue length $\sum_{i}q_{i}$ in the heavy traffic limit (Stolyar, 2004).
6.3. Framework for bounding $V_{\pi}(s)$
@@
6.4. Bounding $V_{MW}(s)$ for MaxWeight
Now, we use the framework established in Section 6.3 to bound the relative value function for the MaxWeight policy.
@@ 6.4. Bounding $V_{MW}(s)$ for MaxWeight
$c_{1}$ $=2\epsilon\min_{i}\lambda_{i},\qquad c_{2}=\ell^{2}n$
@@
This result is proven in (Srikant and Ying, 2013, Theorem 4.2.4), for the classic switch model. However, the proof applies unchanged to any GSSE system. For completeness, we reprove the result in the GSSE setting in Section C.1. ∎
@@ 6.5. Rapid convergence for NPG on queueing MDPs
Theorem 3 ().
For any GSSE queueing system, with the objective of minimizing mean queue length starting with the initial policy $\pi_{0}=$MaxWeight, the NPG algorithm with learning rate parameterization $\beta_{s}$ given in Theorem 1 achieves the convergence rate $E[q_{\pi_{T}}]\geq E[q_{*}]-\frac{c_{*}}{\sqrt{T}}$ for a constant $c_{*}$ given in the proof of Theorem 1.
@@
To prove this result, we will apply Theorem 1. To do so, we will verify Assumptions 1 and 2 and prove the MaxWeight achieves the required bound (4) on the relative value function.
@@
Thus, Theorem 1 applies to any GSSE queueing MDP, given the MaxWeight initial policy. ∎
This result can be generalized to other reward functions, including a weighted queue length reward, as well as the $\alpha$-moment reward function $-(\sum_{i}q_{i})^{\alpha}$, for any $\alpha\geq 1$. For these generalizations of the reward function, we can use weighted-MaxWeight and $\alpha$-MaxWeight as the initial policies. These policies replace the $q_{i}$ term in the MaxWeight definition (6) with the weighted queue length or $\alpha$th power of queue length. The proof that weighted-MaxWeight has the desired relative value property (4) is essentially immediate, and for $\alpha$-MaxWeight one can use the Lyapunov function $f_{\alpha}(q)=\sum_{i}q_{i}^{\alpha+1}$. Assumptions 2 and 1 are straightforward to verify for the weighted queue length setting, and Assumption 1(a-c) and Assumption 2 are straightforward to verify for the $\alpha$th moment reward function. Assumption 1(d) is slightly more involved, but we verify that it holds in Section C.2.
@@ 8. Conclusion
We give the first proof of convergence for the Natural Policy Gradient algorithm in the infinite-state average-reward setting. In particular, we demonstrate that by setting the learning rate function $\beta_{s}$ appropriately, we can guarantee an $O(1/\sqrt{T})$ convergence rate to the optimal policy, as long as the initial policy has a well-behaved relative value function. Moreover, we demonstrate that in infinite-state average-reward MDPs arising out of queueing theory, the MaxWeight policy satisfies our requirement on the initial policy, allowing our results on the NPG algorithm to apply.
@@ Appendix A Background Lemmas
$Q_{k}(s,\pi_{k+1})$ $=\sum_{a}\pi_{k+1}(a\mid s)Q_{k}(s,a)=\sum_{a}\pi_{k+1}(a\mid s)\frac{1}{\log\beta_{s}}\log\left(\frac{Z_{s,k}\pi_{k+1}(a\mid s)}{\pi_{k}(a\mid s)}\right)$
@@
Note that the left-hand term in (8) is simply the $KL$ divergence $D_{KL}(\pi_{k+1}(\cdot|s)||\pi_{k}(\cdot|s))$. As a result, it is positive. Thus,
$Q_{k}(s,\pi_{k+1})$ $\geq\frac{1}{\log\beta_{s}}\sum_{a}\pi_{k+1}(a\mid s)\log Z_{s,k}=\frac{1}{\log\beta_{s}}\log Z_{s,k}=\frac{1}{\log\beta_{s}}\log\sum_{a^{\prime}}\pi_{k}(a^{\prime}\mid s)\beta_{s}^{Q_{k}(s,a^{\prime})}$
@@
$Q_{k}(s,\pi_{k+1})$ $\geq\frac{1}{\log\beta_{s}}\sum_{a^{\prime}}\pi_{k}(a^{\prime}\mid s)Q_{k}(s,a^{\prime})\log\beta_{s}=\sum_{a^{\prime}}\pi_{k}(a^{\prime}\mid s)Q_{k}(s,a^{\prime})=Q_{k}(s,\pi_{k})=V_{k}(s).\qed$
@@ Appendix B Bounding relative value
Lemma 7 states that $c_{5}V_{0}(s)-c_{6}\leq V_{k}(s)\leq c_{7}$. As a result,
$c_{5}E_{s^{\prime\prime}\sim P(s,a^{\prime})}[V_{0}(s^{\prime\prime})]-c_{6}\leq E_{s^{\prime}\sim P(s,a)}[V_{k}(s^{\prime})]\leq c_{7}.$
To relate $V_{0}(s^{\prime\prime})$ to $\hat{r}_{\max}(s)$, we will combine (9), our assumption on $V_{0}(s)$, with Assumption 1(d):
$V_{0}(s^{\prime\prime})$ $\geq-c_{0}\hat{r}_{\max}(s^{\prime\prime})^{2}-c_{1}\geq-c_{0}(R_{3}\hat{r}_{\max}(s)-R_{4})^{2}-c_{1}=-c_{0}(R_{3}|\hat{r}_{\max}(s)|+R_{4})^{2}-c_{1}$
@@
$Q_{k}(s,a)-Q_{k}(s,a^{\prime})$ $\leq c_{0}c_{5}(R_{3}|\hat{r}_{\max}(s)|+R_{4})^{2}+c_{1}c_{5}+c_{6}+c_{6}+R_{1}|\hat{r}_{\max}(s)|^{2}+R_{2}$
$=(c_{0}c_{5}R_{3}^{2}+R_{1})|\hat{r}_{\max}(s)|^{2}+2c_{0}c_{5}R_{3}R_{4}|\hat{r}_{\max}(s)|+c_{0}c_{5}R_{4}^{2}+c_{1}c_{5}+c_{6}+c_{6}+R_{2}$
$=c_{2}|\hat{r}_{\max}(s)|^{2}+c_{3}|\hat{r}_{\max}(s)|+c_{4}$
where $\text{where }c_{2}$ $:=c_{0}c_{5}R_{3}^{2}+R_{1},c_{3}:=2c_{0}c_{5}R_{3}R_{4},\text{ and }c_{4}:=c_{0}c_{5}R_{4}^{2}+c_{1}c_{5}+c_{6}+c_{6}+R_{2}.$
Note that $c_{2},c_{3},$ and $c_{4}$ are all nonnegative. ∎

### v2

@@ 5.5. Step 3: Bounding relative value $V_{k}(s)$ for all states $s$
$\forall s,k,\quad$ $V_{k}(s)\geq\min(V_{0}(s)\frac{J_{*}-z}{J_{0}-z},V_{0}(s))-\tau^{bound}_{J_{0},z}\frac{(J_{*}-J_{0})(c_{\max}-z)(J_{*}-z)}{(J_{0}-z)^{2}}$
@@
We can upper bound each of these quantities. Note that $\bar{r}_{k}^{out}(s)\leq z$, and $\bar{r}_{k}^{in}(s)\leq c_{\max}$. Note that $\tau_{k}^{in}(s)\leq\tau^{bound}_{J_{0},z}$, by Lemma 7. Together, we can upper bound $V_{k}(s)$ relative to $\tau_{k}(s)$, or equivalently upper bound $\tau_{k}(s)$ relative to $V_{k}(s)$:
@@
Recall that by Lemma 3, $J_{\pi}\geq J_{0}>\pi$, so $-J_{k}+z<0$, so the direction of the inequality flips. Here we use Lemma 5, which states that all NPG iterates are stable.
@@
$V_{k+1}(s)$ $\geq V_{k}(s)\frac{J_{k+1}-z}{J_{k}-z}-\tau^{bound}_{J_{0},z}(J_{k+1}-J_{k})\frac{(c_{\max}-z)}{J_{k}-z}$
@@
$V_{k}(s)$ $\geq\min(V_{0}(s)\frac{J_{*}-z}{J_{0}-z},V_{0}(s))-\tau^{bound}_{J_{0},z}\frac{(J_{*}-J_{0})(c_{\max}-z)(J_{*}-z)}{(J_{0}-z)^{2}}$
@@
5.6. Step 4: Proof of general result
@@
Note that for any specific state $s$, $\sum_{k=1}^{T}Q_{k}(s,\pi_{*})-V_{k}(s)$ is exactly the difference in total reward between a specific fixed policy $\pi_{*}(a\mid s)$ and the weighted majority policy $\pi_{k}(a\mid s)$ of the expert advice problem with reward function at time step $k$ of $Q_{k}(s,a)$.
@@
$J_{*}-J_{T}$ $\leq\frac{c_{*}}{\sqrt{T}}.$
5.7. Bounding Relative Value $V_{MW}(s)$ for MaxWeight
@@
Now, we use the framework established in Lemma 10 to bound the relative value function for the MaxWeight policy.
@@
$c_{1}$ $=2\epsilon\min_{i}\lambda_{i},\qquad c_{2}=\ell^{2}n.$
@@
This result is proven in (Srikant and Ying, 2013, Theorem 4.2.4), for the classic $n\times n$ switch model. However, the proof applies unchanged to any GSSE system. For completeness, we reprove the result in the GSSE setting in Section C.1. ∎
@@ 5.8. Rapid convergence for NPG on queueing MDPs
Theorem 1 0.
For any GSSE queueing system, with the objective of minimizing mean queue length starting with the initial policy $\pi_{0}=$MaxWeight, the NPG algorithm with learning rate parameterization $\beta_{s}$ given in Theorem 2 achieves the convergence rate $E[|q_{\pi_{T}}|_{1}]\leq E[|q_{*}|_{1}]+\frac{c_{*}}{\sqrt{T}}$ for a constant $c_{*}$ given in the proof of Theorem 2.
This result generalizes to the $\alpha$th moment of queue length.
@@
To prove this result, we will apply Theorem 2, with reward equal to negative total queue length $-|q|_{1}$. To do so, we will verify Assumptions 1 and 2 and prove the MaxWeight achieves the required bound (9) on the relative value function.
@@
Thus, Theorem 2 applies to any GSSE queueing MDP, given the MaxWeight initial policy.
This result can be generalized to other reward functions, including a weighted queue length reward, as well as the $\alpha$-moment reward function $-(\sum_{i}q_{i})^{\alpha}$, for any $\alpha\geq 1$. For these generalizations of the reward function, we can use weighted-MaxWeight and $\alpha$-MaxWeight as the initial policies. These policies replace the $q_{i}$ term in the MaxWeight definition (3) with the weighted queue length or $\alpha$th power of queue length. The proof that weighted-MaxWeight has the desired relative value property (9) is essentially immediate, and for $\alpha$-MaxWeight one can use the Lyapunov function $f_{\alpha}(q)=\sum_{i}q_{i}^{\alpha+1}$. Assumptions 2 and 1 are straightforward to verify for the weighted queue length setting, and Assumption 1(a-c) and Assumption 2 are straightforward to verify for the $\alpha$th moment reward function. Assumption 1(d) is slightly more involved, but we verify that it holds in Section C.2. ∎
@@ 7. Conclusion
We give the first proof of convergence for the Natural Policy Gradient algorithm in the infinite-state average-reward queueing setting. In particular, we demonstrate that by setting the learning rate function $\beta_{s}$ appropriately, we can guarantee an $O(1/\sqrt{T})$ convergence rate to the optimal policy, as long as the initial policy has a well-behaved relative value function. Moreover, we demonstrate that in infinite-state average-reward MDPs arising out of queueing theory, the MaxWeight policy satisfies our requirement on the initial policy, allowing our results on the NPG algorithm to apply.
@@ Appendix A Background Lemmas
$Q_{k}(s,\pi_{k+1})$ $=\sum_{a}\pi_{k+1}(a\mid s)Q_{k}(s,a)$
(12) $=\sum_{a}\pi_{k+1}(a\mid s)\frac{1}{\log\beta_{s}}\log\left(\frac{Z_{s,k}\pi_{k+1}(a\mid s)}{\pi_{k}(a\mid s)}\right)$
@@
Note that the left-hand term in (13) is simply the $KL$ divergence $D_{KL}(\pi_{k+1}(\cdot\mid s)||\pi_{k}(\cdot\mid s))$. As a result, it is positive. Thus,
$Q_{k}(s,\pi_{k+1})$ $\geq\frac{1}{\log\beta_{s}}\sum_{a}\pi_{k+1}(a\mid s)\log Z_{s,k}$
$=\frac{1}{\log\beta_{s}}\log Z_{s,k}=\frac{1}{\log\beta_{s}}\log\sum_{a^{\prime}}\pi_{k}(a^{\prime}\mid s)\beta_{s}^{Q_{k}(s,a^{\prime})}$
@@
$Q_{k}(s,\pi_{k+1})$ $\geq\frac{1}{\log\beta_{s}}\sum_{a^{\prime}}\pi_{k}(a^{\prime}\mid s)Q_{k}(s,a^{\prime})\log\beta_{s}$
$=\sum_{a^{\prime}}\pi_{k}(a^{\prime}\mid s)Q_{k}(s,a^{\prime})=Q_{k}(s,\pi_{k})=V_{k}(s).$
@@ Appendix B Bounding relative value
Lemma 8 states that $\min(c_{5}V_{0}(s),V_{0}(s))-c_{6}\leq V_{k}(s)\leq c_{7}$. Our lower bound on $V_{0}(s)$, (14), is negative, so we can simplify the bound to $c_{5}\widetilde{V_{0}}(s)-c_{6}\leq V_{k}(s)\leq c_{7}$, where $\widetilde{V_{0}}$ is our lower bound on $V_{0}$. As a result,
$c_{5}E_{s^{\prime\prime}\sim P(s,a^{\prime})}[\widetilde{V_{0}}(s^{\prime\prime})]-c_{6}$ $\leq E_{s^{\prime\prime}\sim P(s,a^{\prime})}[V_{k}(s^{\prime\prime})]$
$E_{s^{\prime}\sim P(s,a)}[V_{k}(s^{\prime})]$ $\leq c_{7}.$
To relate $\widetilde{V_{0}}(s^{\prime\prime})$ to $\hat{r}_{\max}(s)$, we will combine (14), our assumption on $V_{0}(s)$, with Assumption 1(d):
$\widetilde{V_{0}}(s^{\prime\prime})$ $\geq-c_{0}\hat{r}_{\max}(s^{\prime\prime})^{2}-c_{1}\geq-c_{0}(R_{3}\hat{r}_{\max}(s)-R_{4})^{2}-c_{1}$
$=-c_{0}(R_{3}|\hat{r}_{\max}(s)|+R_{4})^{2}-c_{1}$
@@
$Q_{k}(s,a)-Q_{k}(s,a^{\prime})$
$\leq c_{0}c_{5}(R_{3}|\hat{r}_{\max}(s)|+R_{4})^{2}+c_{1}c_{5}+c_{6}+c_{7}+R_{1}|\hat{r}_{\max}(s)|^{2}+R_{2}$
$=(c_{0}c_{5}R_{3}^{2}+R_{1})|\hat{r}_{\max}(s)|^{2}+2c_{0}c_{5}R_{3}R_{4}|\hat{r}_{\max}(s)|+c_{0}c_{5}R_{4}^{2}+c_{1}c_{5}+c_{6}+c_{7}+R_{2}$
$=c_{2}|\hat{r}_{\max}(s)|^{2}+c_{3}|\hat{r}_{\max}(s)|+c_{4},$
where $c_{2}:=c_{0}c_{5}R_{3}^{2}+R_{1},c_{3}:=2c_{0}c_{5}R_{3}R_{4}$, and $c_{4}:=c_{0}c_{5}R_{4}^{2}+c_{1}c_{5}+c_{6}+c_{7}+R_{2}$. Note that $c_{2},c_{3},$ and $c_{4}$ are all nonnegative. ∎